The Arun ice cream’s lowers the price of its vanilla ice cream from Rs.55/- per kg to Rs.40/ kg. Vanilla ice cream sales increases by 20 %. The company notices that the sales of chocolate syrup increased by 10 %.

 

a. What is the price elasticity co-efficient of vanilla ice cream?

b. Why have the sales of chocolate syrup increased and how would you measure the effect?

c. Overall, do you think that the new price policy was beneficial for the Arun ice cream?



1
Expert's answer
2022-01-13T09:09:35-0500

"A)\ PE=\frac{\Delta inVQ\ Demanded}{\Delta in V price}\\
\frac{-15}{55}\times 100\%=-27.28\\
\frac{20\%}{-27.28}=-\ .733"

Less than one demand is price inelastic.

B) they are complimentary goods ..they are jointly demanded hence the increase in demand of chocolate syrup. The effect is measured using cross price elasticity of demand.

C) Yes. Since overall aggregate demand increased .